# Refined asymptotics for the Cauchy problem for the fast $p$-Laplace evolution equation

### v2

@@ 3.4 Comparison principles
By the comparison principle, we mean that, in some sense, ordered data generate ordered solutions at all times. Such results for solutions to (CPLE) seem to be well known by experts in the field, cf. [46, Sections 3 and 4] or [11, Section 4.5]. Nonetheless, we could not find references with complete proofs in the case of the Cauchy problem within the whole range $21<p<$ 2. One of the main difficulties in the proofs of comparison principles is that, at least when $p<p_{c}$, in general solutions are not integrable. Thus, a priori the quantity $(u_{1}-u_{2})_{+}$ (where $u_{1}$ and $u_{2}$ are two solutions to (CPLE)) cannot be used as a test function. However, in our case when solutions are regular and bounded, one is equipped with two comparison principles. We decided to include them with the sketches of the proofs for completeness.
@@
It can be proven via the construction inspired by [42, Chapter 3] where the authors constructed solutions to (CPLE) by approximation by solutions to the Dirichlet problem. The comparison principle for the Dirichlet problem goes back to [68]. For a more recent proof we refer to [41, Chapter 7, Corollary 1.1], cf. also references therein. We notice that for the Dirichlet problem, there is no restriction on $p>1$. We have $u_{1}$ and $u_{2}$ obtained by approximation with solutions to the Dirichlet problem with initial data $u_{1,0}\phi_{R}$ and $u_{2,0}\phi_{R}$ (where $\phi_{R}$ is a cut-off function supported in the ball of radius $R>0$). The approximate solutions are ordered on $B_{2R}$. This relation holds in the limit $R\rightarrow\infty$.
One is not deprived from comparison results in the class of less regular solutions. In fact the following $\mathrm{L}^{1}_{\rm loc}$-comparison principle holds: let $u_{1}$ and $u_{2}$ be two solutions (from the class the class $\Sigma^{\star}$, see [42, Chapter II]):
@@
The main advantage of this local comparison is that it avoids using global integrability, even if it assumes an order in $\mathrm{L}^{1}_{\rm loc}(\mathbb{R}^{N})$. It might be proven by following the lines of [42, Proposition II.3.1 and Theorem II.1.1]. Indeed, a careful inspection of the proof of [42, Proposition II.3.1] shows that under the assumption of (3.4) one gets that, for all $T>0$, $(u_{1}(t,x)-u_{2}(t,x))_{+}\in\mathrm{L}^{\infty}\left((0,T);\mathrm{L}^{q}_{\rm loc}(\mathbb{R}^{N})\right)$ for all $q\in\left[1,\infty\right)$, and there exists $C=C(N,p,q)>0$ such that for all $R>0$, $t\in(0,T)$, and $\sigma\in(0,1)$ one has
@@
Indeed, we also notice that, while originally inequality (3.5) is stated for the absolute value of the difference (i.e. $|u_{1}-u_{2}|$), its proof is done for the positive part of $u_{1}-u_{2}$. Once (3.5) is obtained, using the same argument as in the proof of [42, Theorem II.1.1, p. 257] one obtains that, for every $q>1$, all $t\in\left(0,T\right)$, and $C=C(N,p,q)$ independent of $R>0$:
@@
By choosing $q$ so large such that $qp/(2-p)>N$, one obtains in the limit $R\rightarrow 0$ that $\int_{\mathbb{R}^{N}}(u_{1}(t,x)-u_{2}(t,x))_{+}^{q}\,{\rm d}x=0$ for all $t>0$. Consequently, $u_{1}(t,x)\leq u(t_{2},x)$ for all $t>0$ and $x\in\mathbb{R}^{N}$. We acknowledge that, despite the comparison principle has not been stated in the form (3.4) in [42], it has been used in this form in [42, Proposition III.7.1]. Hence, we do not claim any originality for the above result.
@@ 4.3 Convergence in inner cylinders
$\begin{split}t^{N\beta}\,\|u(t,\cdot)-\mathcal{B}_{M}(t,\cdot)\|_{\mathrm{L}^{\infty}({\mathbb{R}^{N}})}&\leq C_{N}\,t^{N\beta}\,\|\nabla u(t,\cdot)-\nabla\mathcal{B}_{M}(t,\cdot)\|_{\mathrm{L}^{\infty}({\mathbb{R}^{N}})}^{\frac{N}{N+1}}\,\|u(t,\cdot)-\mathcal{B}_{M}(t,\cdot)\|_{\mathrm{L}^{1}({\mathbb{R}^{N}})}^{\frac{1}{N+1}}\\ &\leq C_{N}\,t^{N\beta}\,c_{1}^{\frac{N}{N-1}}\left(2\,Mt^{-(N+1)\beta}\right)^{\frac{N}{N+1}}\left(\widetilde{K}\,t^{-\nu}\right)^{\frac{1}{N+1}}\\ &\leq C_{\star}\,t^{N\beta-\frac{N}{N+1}(N+1)\beta-\frac{\nu}{N+1}}\leq C_{\star}\,t^{-\frac{\nu}{N+1}}\,,\end{split}$
@@ 4.4 Proof of convergence in relative error under a priori convergence in Lebesgue’s space
[…]
@@ 5.1 Relation between the relative entropy and the Fisher information along the flow
It will be clear in a few lines that the above quantities are well defined. In what follows, we would like to test equation (R-CPLE) against $\gamma(v^{\gamma-1}-V_{D}^{\gamma-1})\phi_{R}$. This is an admissible test function. Indeed, the assumption (A0) on ${{v_{0}}}$ implies (by the maximum principle) that $v(\tau)$ satisfies (A0) for all $\tau>0$. Moreover, we know that for a fixed $R>0$ function $y\mapsto v^{\gamma-1}\phi_{R}$ is $C^{1}$. Actually more is known: the function $y\mapsto v(\tau,y)$ is $C^{1,\alpha}$ locally in space for some $\alpha=\alpha(N,p)$, i.e., $\|\nabla v\|_{C^{0,\alpha}(B_{R})}$ is finite for any $R>0$, see [42, Theorem III.8.1]. Therefore, by a standard approximation procedure, we can test (R-CPLE) against $\gamma(v^{\gamma-1}-V_{D}^{\gamma-1})\phi_{R}$ to get
@@
${\mathcal{E}}_{\phi_{R}}[v(\tau)|{V_{D}}]$ $R\to\infty]{}{\mathcal{E}}[v(\tau)|{V_{D}}]\,,$ (5.5)
$\int_{\tau_{0}}^{\tau_{0}+h}{\mathcal{I}}_{\phi_{R}}[v(\tau)|{V_{D}}]\,{\rm d}\tau$ $R\to\infty]{}\int_{\tau_{0}}^{\tau_{0}+h}{\mathcal{I}}[v(\tau)|{V_{D}}]\,{\rm d}\tau\,,$ (5.6)
$\int_{\tau_{0}}^{\tau_{0}+h}\big{|}\mathcal{R}_{\phi_{R}}[v(\tau)|{V_{D}}]\big{|}$ $R\to\infty]{}0\,.$ (5.7)
@@
$\begin{split}|F_{R}(y)|&\leq\tfrac{C_{1}}{R}\,|V_{D}(y)|^{\gamma-1}\left(|\nabla v|^{p-1}+|y|\,|vV_{D}(y)|\right)\leq\tfrac{C_{2}}{R}\,|vV_{D}(y)|^{\gamma-1}\,|y|\left({|y|^{-\frac{p}{2-p}}}+|V_{D}(y)|\right)\\ &\leq C_{3}\,|V_{D}(y)|^{\gamma-1}\left({|y|^{-\frac{p}{2-p}}}+|V_{D}(y)|\right)\leq C_{4}\,|V_{D}(y)|^{\gamma}\,,\end{split}$
@@
On the right-hand side above, the second integral converges to 00, as $R\rightarrow\infty$, since $v(\tau,\cdot)$ is an integrable function and $|A_{R}|\subset\{|y|>R\}$. It only remains to justify the convergence of the first integral. By inequality (5.2), we have that for large enough $|y|$ and large enough $\tau>0$ it holds $|\nabla v(\tau,y)|^{p-1}\leq c_{5}\,|y|^{-\frac{2(p-1)}{2-p}}$ for a constant $c_{5}>0$ independent of $y$ (a similar estimate holds also for small $\tau$ but with a constant $c_{5}$ depending on $\tau$). Therefore, we are left with
@@ 5.2 Linearised entropy functional and Fisher information
where we used the fact that $V_{D}^{\gamma-2}(y)\,|y|^{-\frac{p}{(2-p)(p-1)}}\leq C(D,N,p)$ for $|y|\geq 1$ where $C(D,N,p)>0$ depends only on $D$, $N$ and $p$. Combining the two cases $|y|\leq 1$ and $|y|\geq 1$ together, we find that (5.14) holds for all $y\in{\mathbb{R}^{N}}$. Then (5.14) together with (5.13) implies (5.12) with the constant $C_{\varepsilon}=2^{p-2}\,C({{v_{0}}},p,N,\varepsilon,D)^{p-2}\,c_{2}(1-\varepsilon)$. ∎
@@ 5.4 Convergence in $\mathrm{L}^{1}$, case $p_{M}<p<2$
Let $N\geq 2$, $p_{M}<p<p_{2}$, $0\leq{{v_{0}}}\in\mathrm{L}^{1}({\mathbb{R}^{N}})\cap\mathcal{X}_{p}$, be such that $\int_{{\mathbb{R}^{N}}}{{v_{0}}}(y)\,{\rm d}y=M_{\star}$. Assume that $v$ is a weak solution to (R-CPLE) with ${{v_{0}}}$ as initial datum. If ${{v_{0}}}$ satisfies (A1) for some ${\varepsilon}\in(0,1)$ and $D=1$, then there exists $\tau_{\varepsilon}>0$,$c=c({{v_{0}}},p,N,{\varepsilon})$, and $\vartheta=\vartheta(p,N,{\varepsilon})>0$ such that for any $\tau>\tau_{\varepsilon}$ it holds
@@ 5.6 Proof of Theorem 1
$\begin{split}\|u(t,\cdot)-\mathcal{B}_{M_{\star}}(t,\cdot)\|_{\mathrm{L}^{1}({\mathbb{R}^{N}})}&\leq\|u(t,x)-\mathcal{B}_{M_{\star}}(t+T,\cdot)\|_{\mathrm{L}^{1}({\mathbb{R}^{N}})}+\|\mathcal{B}_{M_{\star}}(t+T,\cdot)-\mathcal{B}_{M_{\star}}(t,\cdot)\|_{\mathrm{L}^{1}({\mathbb{R}^{N}})}\\ &\leq\tfrac{\widetilde{K}}{t^{\nu}}+\tfrac{C(T)}{t}\leq\tfrac{\widetilde{K}+C(T)}{t^{\nu}}\,,\end{split}$
@@
On the other hand, inequality (5.30) can be proven by considering the relative error between the two solutions $\mathcal{B}_{M_{\star}}(t+T,\cdot)$ and $\mathcal{B}_{M_{\star}}(t,\cdot)$. Indeed, for any fixed $t>0$, the supremum in $|x|$ of the quotient $\mathcal{B}_{M_{\star}}(t+T,\cdot)/\mathcal{B}_{M_{\star}}(t,\cdot)$ is attained either at 00 or at $\infty$ (one can prove this through a simple, however lengthy, computation). From this observation one finds that
@@ 6.1 Information on related radial classical and weighted Fast Diffusion Equations
Below we present a radial equivalence due to [50, Theorem 1.2]. We point out, however, that we use a particular version of the result stated in [50]. The original result is indeed much stronger and valid also for sign-changing solutions. Notice also that we consider a slightly different equation, so the constant $\mathcal{D}$ defined below differs slightly comparing with the one of [50, Theorem 1.2]. In what follows we shall denote by $r=|x|$ in the (6.1) case and by ${\varrho}=|x|$ the coordinates for the (6.2) equation.
Theorem 12 ([50, Theorem 1.2]).
@@
$-\partial_{r}u(t,r)={\mathcal{D}}\,{\varrho}^{\frac{2}{m+1}}\,\Phi(t,{\varrho})\,,\quad\mathcal{D}=\left(\frac{2m}{m+1}\right)^{\frac{2}{m-1}}\,,$ (6.3)
@@
Let us note that in [50] the authors also analyze the case $20<n<$ 2, however, we shall not use those results. We also remark that, even if the solutions to (CPLE) are at least $C^{1,\alpha}$, a priori we do not know whether $\Phi$ is well-defined at the origin by transformation (6.3). We will address these issues below.
@@
$-\partial_{r}\mathcal{B}_{M}(t,r)={\mathcal{D}}\,\varrho^{\frac{2}{m+1}}\mathfrak{B}_{\mathfrak{C}M}(t,\varrho)\,,$ (6.9)
@@
[…]
@@
$-\partial_{r}\mathcal{B}_{D,T}(t,r)={\mathcal{D}}\,\varrho^{\frac{2}{m+1}}\mathfrak{B}_{\overline{\mathfrak{C}}D,T}(t,\varrho)\,,$ (6.12)
where ¯=¯$\overline{\mathfrak{C}}=\overline{\mathfrak{C}}(N,p)>0$ is given by
¯$\overline{\mathfrak{C}}=\left(\frac{1-m}{2m}\right)^{2\theta+1}\,.$ (6.13)
@@
Suppose that $N\geq 2$ , $21\leq p<$ 2, ${\mathcal{D}}$ is as in (6.3), and $n$, $m$, ${\mathfrak{a}}$ are as in (6.4). Let $u$ be a radial solution to (CPLE) with a radial initial datum ${{u_{0}}}\in C^{2}({\mathbb{R}}^{N})$ satisfying one of the following conditions:
@@
$\begin{cases}\partial_{t}\Phi=|x|^{{\mathfrak{a}}}\,\mathrm{div}\left(|x|^{-{\mathfrak{a}}}\,\nabla\Phi^{m}\right)\,\quad\text{for }\ (t,x)\in\left(0,\infty\right)\times\mathbb{R}^{N}\\ \Phi(0,x)=\Phi_{0}(|x|)\,\quad\text{for }\ x\in\mathbb{R}^{N}\,,\end{cases}$ (6.14)
@@
We are in the position to pass to justification of solvability of (6.14). Let us notice that the equation in (6.14) is the same as in (6.2). As explained in (6.6), in the present range of parameters we always have ${\mathfrak{a}}<0$. Solutions for problem (6.14) have been constructed in [16] in the same spirit as in [49, Theorem 2.1]. We stress that in [16, Proposition 7] the initial datum is assumed to be in $\mathrm{L}^{\infty}(\mathbb{R}^{N})$. This assumption can be weakened to merely $\Phi_{0}\in\mathrm{L}^{1}(\mathbb{R}^{N},|x|^{-{\mathfrak{a}}}\,{\rm d}x)$ by a standard approximation procedure as it is done in the proof of [49, Theorem 2.1]. Since $\Phi_{0}\geq 0$ and the comparison principle holds due to [16, Corollary 9], we know that the solution is non-negative. In this range of parameters solutions are bounded since $\Phi_{0}\in\mathrm{L}^{1}_{{\rm loc}}({\mathbb{R}^{N}},|x|^{-{\mathfrak{a}}}\,\,{\rm d}x)$, see [22, Theorem 1.2]. It is also known that solutions are at least $C^{\alpha}$-regular close to the origin (see [22, Theorem 1.8]) and $C^{\infty}$-smooth outside of the origin. This has been already remarked in [16, Lemma 11]. See also [41, Section 21.5.3] where the authors affirm that local analyticity in space and, at least, Lipschitz continuity in time holds for solutions to a general equation of the form (6.14). This considerations prove that the solution $\Phi$ to (6.14) exists and it has the wanted properties. It only remains to verify that a radial solution $u$ to (CPLE) is related to the Cauchy problem (6.14) through the transformation (6.3). Despite this seems obvious, it is not since we need to find a relevant relation between initial data. Let us consider the auxiliary function
@@
where $C(t,\tau)\lesssim t^{-n\,\theta}$ for any $t>2\tau$ with $\theta$ as in (6.7). The exponent $12/(2-p)>$ 1 since $12>p>$ 1, so we deduce from the above inequality that the function $\overline{u}$ is well defined. Furthermore, we have that $\overline{u}(t)\in\mathrm{L}^{\infty}(0,\infty)$ for any $t>0$ (indeed, $\tau_{0}$ is chosen arbitrarily). Let us now investigate the regularity of $\Phi$. The validity of the inequality (6.17) allows us to use the regularity information resulting from the proof of [16, Lemma 11]. By those results we have that $\Phi\in C^{\infty}(0,\infty)^{2}$, and, for any $\tau>0$, $\varepsilon>0$ and $k>0$ there exist $C_{1}=C_{1}(t,\tau,\varepsilon)>0$ and $C_{2}=C_{2}(t,\tau,\varepsilon,k)$ such that
@@
We briefly comment on the literature for the above problem in Remark 6.2. Let us continue with the rest of the proof. By using the result of [42, Theorem III.8.1], we know that the function $(t,x)\mapsto\nabla u(t,x)\in C^{\alpha}_{{\rm loc}}((0,\infty)\times\mathbb{R}^{N})$, which is enough to guarantee that $\partial_{r}{{u}}(t,0)={0}$ for all $t>0$. We conclude therefore that $u$ also solves problem (6.20). Since ${{u_{0}}}=\overline{u}_{0}$ we would like to conclude that $u=\overline{u}$ by using the uniqueness result for (6.20). This would be enough to conclude the proof, since by the construction we will have that $\Phi$ and $u$ satisfy the relation (6.3). However, in order to apply the uniqueness result of [42, Theorem II.1] we need to ensure that
@@
where $\mathfrak{h}=\mathfrak{h}(N,p,t)$ is independent of $\overline{u}$. We notice that, as observed in [42, p. 45], solutions to problem (CPLE) satisfy (6.21) by the construction. In order to prove (6.21) for solutions to (6.20) we shall use a modification of a trick due to Bénilan and Crandall [9] provided in [42, Lemma III.3.4]. By the comparison principle proven in [16], the uniqueness for (6.14) is guaranteed. Let us consider $\Psi^{\lambda}$ being the unique solution to (6.14) with initial datum

### v3

@@ 3.4 Comparison principles
By the comparison principle, we mean that, in some sense, ordered data generate ordered solutions at all times. Such results for solutions to (CPLE) seem to be well known by experts in the field, cf. [47, Sections 3 and 4] or [11, Section 4.5]. Nonetheless, we could not find references with complete proofs in the case of the Cauchy problem within the whole range $1<p<2$. One of the main difficulties in the proofs of comparison principles is that, at least when $p<p_{c}$, in general solutions are not integrable. Thus, a priori the quantity $(u_{1}-u_{2})_{+}$ (where $u_{1}$ and $u_{2}$ are two solutions to (CPLE)) cannot be used as a test function. However, in our case when solutions are regular and bounded, one is equipped with two comparison principles. We decided to include them with the sketches of the proofs for completeness.
@@
It can be proven via the construction inspired by [43, Chapter 3] where the authors constructed solutions to (CPLE) by approximation by solutions to the Dirichlet problem. The comparison principle for the Dirichlet problem goes back to [69]. For a more recent proof we refer to [42, Chapter 7, Corollary 1.1], cf. also references therein. We notice that for the Dirichlet problem, there is no restriction on $p>1$. We have $u_{1}$ and $u_{2}$ obtained by approximation with solutions to the Dirichlet problem with initial data $u_{1,0}\phi_{R}$ and $u_{2,0}\phi_{R}$ (where $\phi_{R}$ is a cut-off function supported in the ball of radius $R>0$). The approximate solutions are ordered on $B_{2R}$. This relation holds in the limit $R\rightarrow\infty$.
One is not deprived from comparison results in the class of less regular solutions. In fact the following $\mathrm{L}^{1}_{\rm loc}$-comparison principle holds: let $u_{1}$ and $u_{2}$ be two solutions (from the class the class $\Sigma^{\star}$, see [43, Chapter II]):
@@
The main advantage of this local comparison is that it avoids using global integrability, even if it assumes an order in $\mathrm{L}^{1}_{\rm loc}(\mathbb{R}^{N})$. It might be proven by following the lines of [43, Proposition II.3.1 and Theorem II.1.1]. Indeed, a careful inspection of the proof of [43, Proposition II.3.1] shows that under the assumption of (3.4) one gets that, for all $T>0$, $(u_{1}(t,x)-u_{2}(t,x))_{+}\in\mathrm{L}^{\infty}\left((0,T);\mathrm{L}^{q}_{\rm loc}(\mathbb{R}^{N})\right)$ for all $q\in\left[1,\infty\right)$, and there exists $C=C(N,p,q)>0$ such that for all $R>0$, $t\in(0,T)$, and $\sigma\in(0,1)$ one has
@@
Indeed, we also notice that, while originally inequality (3.5) is stated for the absolute value of the difference (i.e. $|u_{1}-u_{2}|$), its proof is done for the positive part of $u_{1}-u_{2}$. Once (3.5) is obtained, using the same argument as in the proof of [43, Theorem II.1.1, p. 257] one obtains that, for every $q>1$, all $t\in\left(0,T\right)$, and $C=C(N,p,q)$ independent of $R>0$:
@@
By choosing $q$ so large such that qp/(2p)>Nqp/(2-p)>N, one obtains in the limit $R\rightarrow 0$ that $\int_{\mathbb{R}^{N}}(u_{1}(t,x)-u_{2}(t,x))_{+}^{q}\,{\rm d}x=0$ for all $t>0$. Consequently, $u_{1}(t,x)\leq u(t_{2},x)$ for all $t>0$ and $x\in\mathbb{R}^{N}$. We acknowledge that, despite the comparison principle has not been stated in the form (3.4) in [43], it has been used in this form in [43, Proposition III.7.1]. Hence, we do not claim any originality for the above result.
@@ 4.3 Convergence in inner cylinders
[…]
@@ 4.4 Proof of convergence in relative error under a priori convergence in Lebesgue’s space
$\left\|\frac{u(t,\cdot)}{\mathcal{B}_{M}(t,\cdot)}-1\right\|_{\mathrm{L}^{\infty}({\mathbb{R}^{N}})}\leq\varepsilon\leq 2^{-m}\leq 2\,2^{-(m+1)}\leq 2\,C_{p}^{\frac{(2-p)\nu}{N+1}}\,t^{-\frac{(2-p)\nu}{N+1}}\,.$
@@ 5.1 Relation between the relative entropy and the Fisher information along the flow
It will be clear in a few lines that the above quantities are well defined. In what follows, we would like to test equation (R-CPLE) against $\gamma(v^{\gamma-1}-V_{D}^{\gamma-1})\phi_{R}$. This is an admissible test function. Indeed, the assumption (A0) on $v_{0}$}} implies (by the maximum principle) that $v(\tau)$ satisfies (A0) for all $\tau>0$. Moreover, we know that for a fixed $R>0$ function $y\mapsto v^{\gamma-1}\phi_{R}$ is $C^{1}$. Actually more is known: the function $y\mapsto v(\tau,y)$ is $C^{1,\alpha}$ locally in space for some $\alpha=\alpha(N,p)$, i.e., $\|\nabla v\|_{C^{0,\alpha}(B_{R})}$ is finite for any $R>0$, see [43, Theorem III.8.1]. Therefore, by a standard approximation procedure, we can test (R-CPLE) against $\gamma(v^{\gamma-1}-V_{D}^{\gamma-1})\phi_{R}$ to get
@@
${\mathcal{E}}_{\phi_{R}}[v(\tau)|{V_{D}}]$ $\xrightarrow[R\to\infty]{}{\mathcal{E}}[v(\tau)|{V_{D}}]\,$, (5.5)
$\int_{\tau_{0}}^{\tau_{0}+h}{\mathcal{I}}_{\phi_{R}}[v(\tau)|{V_{D}}]\,{\rm d}\tau$ $\xrightarrow[R\to\infty]{}\int_{\tau_{0}}^{\tau_{0}+h}{\mathcal{I}}[v(\tau)|{V_{D}}]\,{\rm d}\tau\,$, (5.6)
$\int_{\tau_{0}}^{\tau_{0}+h}\big|\mathcal{R}_{\phi_{R}}[v(\tau)|{V_{D}}]\big|$ $\xrightarrow[R\to\infty]{}0\,.$ (5.7)
@@
[…]
@@
On the right-hand side above, the second integral converges to 0, as $R\rightarrow\infty$, since $v(\tau,\cdot)$ is an integrable function and $|A_{R}|\subset\{|y|>R\}$. It only remains to justify the convergence of the first integral. By inequality (5.2), we have that for large enough $|y|$ and large enough $\tau>0$ it holds $|\nabla v(\tau,y)|^{p-1}\leq c_{5}\,|y|^{-\frac{2(p-1)}{2-p}}$ for a constant $c_{5}>0$ independent of $y$ (a similar estimate holds also for small $\tau$ but with a constant $c_{5}$ depending on $\tau$). Therefore, we are left with
@@ 5.2 Linearised entropy functional and Fisher information
where we used the fact that $V_{D}^{\gamma-2}(y)\,|y|^{-\frac{p}{(2-p)(p-1)}}\leq C(D,N,p)$ for $|y|\geq 1$ where $C(D,N,p)>0$ depends only on $D$, NN and $p$. Combining the two cases $|y|\leq 1$ and $|y|\geq 1$ together, we find that (5.14) holds for all $y\in{\mathbb{R}^{N}}$. Then (5.14) together with (5.13) implies (5.12) with the constant $C_{\varepsilon}=2^{p-2}\,C({{v_{0}}},p,N,\varepsilon,D)^{p-2}\,c_{2}(1-\varepsilon)$. ∎
@@ 5.4 Convergence in $\mathrm{L}^{1}$, case $p_{M}<p<2$
Let $N\geq 2$, $p_{M}<p<p_{2}$, $0\leq{{v_{0}}}\in\mathrm{L}^{1}({\mathbb{R}^{N}})\cap\mathcal{X}_{p}$, be such that $\int_{{\mathbb{R}^{N}}}{{v_{0}}}(y)\,{\rm d}y=M_{\star}$. Assume that $v$ is a weak solution to (R-CPLE) with $v_{0}$}} as initial datum. If $v_{0}$}} satisfies (A1) for some ${\varepsilon}\in(0,1)$ and $D=1$, then there exists $\tau_{\varepsilon}>0$$c=c({{v_{0}}},p,N,{\varepsilon})$, and $\vartheta=\vartheta(p,N,{\varepsilon})>0$ such that for any $\tau>\tau_{\varepsilon}$ it holds
@@ 5.6 Proof of Theorem 1
[…]
@@
On the other hand, inequality (5.30) can be proven by considering the relative error between the two solutions $\mathcal{B}_{M_{\star}}(t+T,\cdot)$ and $\mathcal{B}_{M_{\star}}(t,\cdot)$. Indeed, for any fixed $t>0$, the supremum in $|x|$ of the quotient $\mathcal{B}_{M_{\star}}(t+T,\cdot)/\mathcal{B}_{M_{\star}}(t,\cdot)$ is attained either at 0 or at $\infty$ (one can prove this through a simple, however lengthy, computation). From this observation one finds that
@@ 6.1 Information on related radial classical and weighted Fast Diffusion Equations
Below we present a radial equivalence due to [51, Theorem 1.2]. We point out, however, that we use a particular version of the result stated in [51]. The original result is indeed much stronger and valid also for sign-changing solutions. Notice also that we consider a slightly different equation, so the constant $\mathcal{D}$ defined below differs slightly comparing with the one of [51, Theorem 1.2]. In what follows we shall denote by $r=|x|$ in the (6.1) case and by ${\varrho}=|x|$ the coordinates for the (6.2) equation.
Theorem 12 ([51, Theorem 1.2]).
@@
$\partial_{r}u(t,r)={\mathcal{D}}\,{\varrho}^{\frac{2}{m+1}}\,\Phi(t,{\varrho})\,,\quad\mathcal{D}=\left(\frac{2m}{m+1}\right)^{\frac{2}{m-1}}\,$, (6.3)
@@
Let us note that in [51] the authors also analyze the case $0<n<2$, however, we shall not use those results. We also remark that, even if the solutions to (CPLE) are at least $C^{1,\alpha}$, a priori we do not know whether $\Phi$ is well-defined at the origin by transformation (6.3). We will address these issues below.
@@
$\partial_{r}\mathcal{B}_{M}(t,r)={\mathcal{D}}\,\varrho^{\frac{2}{m+1}}\mathfrak{B}_{\mathfrak{C}M}(t,\varrho)\,$, (6.9)
@@
$\mathfrak{R}_{T}(t):=\left(\frac{T-t}{|\theta|}\right)_{+}^{\theta}\quad\mbox{if}\,\,0<m<\frac{n-2}{n}\quad\mbox{and}\quad\mathfrak{R}_{T}(t):=\exp\{\mathfrak{l}(t+T)\}\quad\mbox{if}\,\,m=\frac{n-2}{n}\,$, (6.11)
@@
$\partial_{r}\mathcal{B}_{D,T}(t,r)={\mathcal{D}}\,\varrho^{\frac{2}{m+1}}\mathfrak{B}_{\overline{\mathfrak{C}}D,T}(t,\varrho)\,$, (6.12)
where $\overline{\mathfrak{C}}=\overline{\mathfrak{C}}(N,p)>0$ is given by
$\overline{\mathfrak{C}}=\left(\frac{1-m}{2m}\right)^{2\theta+1}\,.$ (6.13)
@@
Suppose that $N\geq 2$ , $1\leq p<2$, $\mathcal{D}$} is as in (6.3), and $n$, $m$, $\mathfrak{a}$} are as in (6.4). Let $u$ be a radial solution to (CPLE) with a radial initial datum ${{u_{0}}}\in C^{2}({\mathbb{R}}^{N})$ satisfying one of the following conditions:
@@
[…]
@@
We are in the position to pass to justification of solvability of (6.14). Let us notice that the equation in (6.14) is the same as in (6.2). As explained in (6.6), in the present range of parameters we always have ${\mathfrak{a}}<0$. Solutions for problem (6.14) have been constructed in [16] in the same spirit as in [50, Theorem 2.1]. We stress that in [16, Proposition 7] the initial datum is assumed to be in $\mathrm{L}^{\infty}(\mathbb{R}^{N})$. This assumption can be weakened to merely $\Phi_{0}\in\mathrm{L}^{1}(\mathbb{R}^{N},|x|^{-{\mathfrak{a}}}\,{\rm d}x)$ by a standard approximation procedure as it is done in the proof of [50, Theorem 2.1]. Since $\Phi_{0}\geq 0$ and the comparison principle holds due to [16, Corollary 9], we know that the solution is non-negative. In this range of parameters solutions are bounded since $\Phi_{0}\in\mathrm{L}^{1}_{{\rm loc}}({\mathbb{R}^{N}},|x|^{-{\mathfrak{a}}}\,\,{\rm d}x)$, see [22, Theorem 1.2]. It is also known that solutions are at least $C^{\alpha}$-regular close to the origin (see [22, Theorem 1.8]) and $C^{\infty}$-smooth outside of the origin. This has been already remarked in [16, Lemma 11]. See also [42, Section 21.5.3] where the authors affirm that local analyticity in space and, at least, Lipschitz continuity in time holds for solutions to a general equation of the form (6.14). This considerations prove that the solution $\Phi$ to (6.14) exists and it has the wanted properties. It only remains to verify that a radial solution $u$ to (CPLE) is related to the Cauchy problem (6.14) through the transformation (6.3). Despite this seems obvious, it is not since we need to find a relevant relation between initial data. Let us consider the auxiliary function
@@
where $C(t,\tau)\lesssim t^{-n\,\theta}$ for any $t>2\tau$ with $\theta$ as in (6.7). The exponent 2/(2p)>12/(2-p)>1 since $2>p>1$, so we deduce from the above inequality that the function $\overline{u}$ is well defined. Furthermore, we have that $\overline{u}(t)\in\mathrm{L}^{\infty}(0,\infty)$ for any $t>0$ (indeed, $\tau_{0}$ is chosen arbitrarily). Let us now investigate the regularity of $\Phi$. The validity of the inequality (6.17) allows us to use the regularity information resulting from the proof of [16, Lemma 11]. By those results we have that $\Phi\in C^{\infty}(0,\infty)^{2}$, and, for any $\tau>0$, $\varepsilon>0$ and $k>0$ there exist $C_{1}=C_{1}(t,\tau,\varepsilon)>0$ and $C_{2}=C_{2}(t,\tau,\varepsilon,k)$ such that
@@
We briefly comment on the literature for the above problem in Remark 6.2. Let us continue with the rest of the proof. By using the result of [43, Theorem III.8.1], we know that the function $(t,x)\mapsto\nabla u(t,x)\in C^{\alpha}_{{\rm loc}}((0,\infty)\times\mathbb{R}^{N})$, which is enough to guarantee that $\partial_{r}{{u}}(t,0)={0}$ for all $t>0$. We conclude therefore that $u$ also solves problem (6.20). Since ${{u_{0}}}=\overline{u}_{0}$ we would like to conclude that $u=\overline{u}$ by using the uniqueness result for (6.20). This would be enough to conclude the proof, since by the construction we will have that $\Phi$ and $u$ satisfy the relation (6.3). However, in order to apply the uniqueness result of [43, Theorem II.1] we need to ensure that
@@
where $\mathfrak{h}=\mathfrak{h}(N,p,t)$ is independent of $\overline{u}$. We notice that, as observed in [43, p. 45], solutions to problem (CPLE) satisfy (6.21) by the construction. In order to prove (6.21) for solutions to (6.20) we shall use a modification of a trick due to Bénilan and Crandall [9] provided in [43, Lemma III.3.4]. By the comparison principle proven in [16], the uniqueness for (6.14) is guaranteed. Let us consider $\Psi^{\lambda}$ being the unique solution to (6.14) with initial datum